OpenStudy (anonymous):

Is it 11 or -11? What is the distance between points (9, –7) and (9, 4) on a coordinate plane?

OpenStudy (anonymous):

1,3 i think

OpenStudy (anonymous):

1,-3

OpenStudy (anonymous):

D= sqrt[(y1-y2)^2+(x1-x2)^2]

OpenStudy (anonymous):

D=sqrt[(-7-(4))^2+(9-9)^2]=sqrt(11)^2=11

OpenStudy (anonymous):

one thing distance always positive

OpenStudy (anonymous):

Oh okie. Thnx! :)